### v2

@@ Abstract
Navigation of wheeled vehicles on uneven terrain necessitates going beyond the 2D approaches for trajectory planning. Specifically, it is essential to incorporate the full $6dof$ variation of vehicle pose and its associated stability cost in the planning process. To this end, most recent works aim to learn a neural network model to predict the vehicle evolution. However, such approaches are data-intensive and fraught with generalization issues.
@@
We improve the state-of-the-art in the following respects. First, we show that our NLS based pose prediction closely matches the output from a high-fidelity physics engine. This result coupled with the fact that we can query gradients of the NLS solver, makes our pose predictor, a differentiable wheel-terrain interaction model. We further leverage this differentiability to efficiently solve the proposed bi-level trajectory optimization problem. Finally, we perform extensive experiments, and comparison with a baseline to showcase the effectiveness of our approach in obtaining smooth, stable trajectories.
@@ I Introduction
Applications like forestry, construction, and search and rescue require wheeled vehicles to navigate over uneven terrains. Thus, for automation in these use cases, we need a trajectory planner that can account for the vehicle tip-over stability over uneven terrains. This in turn requires explicitly predicting the vehicle’s $6dof$ pose on the terrain and the kinematic and stability costs associated with those poses.
@@
Algorithmic Contribution: We jointly view the wheeled vehicle and the underneath terrain through the lens of closed-loop kinematics, typically used for modeling parallel manipulators. This in turn leads to a set of coupled non-linear equations that dictate how the vehicle’s $6dof$ pose depends on the terrain surface geometry. The pose prediction can then be obtained by minimizing the residual of the non-linear equations, which, in our case, is done through a least-squares optimization problem. Importantly, our set-up allows for differentiating through the least squares solver using implicit function theorem to ascertain how the pose of the vehicle will vary if its position on the terrain is perturbed. Thus, in essence, our non-linear least squares (NLS) approach provides a differentiable wheel-terrain interaction model, albeit at the kinematic level.
@@ III-C Tip-over stability criteria
$\boldsymbol{\pi}_{i}$ $=(\mathrm{I}-\mathbf{\hat{e}}_{i}\mathbf{\hat{e}}_{i}^{T})(\mathbf{p}_{oc,i+1}-\mathbf{p}_{og}),$ (11)
$\boldsymbol{\nu}_{i}$ $=(\mathrm{I}-\mathbf{\hat{e}}_{i}\mathbf{\hat{e}}_{i}^{T})\boldsymbol{\nu},$ (12)
where $\mathrm{I}$ is the identity matrix and $\boldsymbol{\nu}$ is the total force acting on the vehicle. The angles between the forces and the tip-over axis normals can be computed as:
@@ IV-A Functional Form for the Terrain Model
$f(x^{j},y^{j})=\sum_{n=1}^{N}a_{n}\cos{(\omega_{1,n}x^{j}+\omega_{2,n}y^{j})}+b_{n}\sin{(\omega_{3,n}x^{j}+\omega_{4,n}y^{j})},$ (20)
@@ IV-B Differentiable Wheel-Terrain Interaction
We can expand (23) using the Euler-angle parametrization for the rotation matrix $\mathbf{R}$. When the resulting expressions are stacked alongside (30), we obtain a set of coupled non-linear equations in the following form:
@@ IV-C Bi-Level Optimization Based Trajectory Planning
where $\mathbf{u}^{*}$ is formed by stacking $\mathbf{u}_{k}^{*}$ at different instants of time. The matrices $\mathbf{A}_{eq},\mathbf{A}$ and the vectors $\mathbf{b}_{eq},\mathbf{b}$ comes from (34b)-(34e).
@@ IV-D Projected gradient descent
The optimal value for the parameter $\boldsymbol{\xi}_{k}$ is obtained through projected gradient descent. The gradient descent update rule is given as
@@
${}^{t+1}\boldsymbol{\xi}=$ $\arg\min_{\boldsymbol{\xi}}\frac{1}{2}\left\lVert{}^{t+1}\overline{\boldsymbol{\xi}}-\boldsymbol{\xi}\right\rVert_{2}^{2},$ (37)
@@ V-A Implementation Details
The trajectory planner comprising of the bi-level optimization (35a)-(35c) and projection (38)-(39) is implemented in Python using the JAX [23] library for GPU-accelerated computing. The matrix W in (9) is constructed from a $10^{th}$ order polynomial. Our simulation pipeline consists of open-loop analyses carried out with stand-alone Python code, and the closed-loop simulations were performed on the Gazebo simulator [24]. The terrains were created using Blender [25] and were converted to the required formats for the Python and Gazebo simulations. All simulations were conducted on a computer with Intel i9 CPU and Nvidea RTX 3080 GPU. For modeling the terrain at each instant of time, we take a patch of radius $7m$ centered at the vehicle position. The number of Fourier frequencies $N$ in the terrain model was taken as $100$ to balance the trade-off between computation time and accuracy.
@@ V-B Validation using the Gazebo simulator
In this section, we analyze the accuracy of our pose predictor by comparing the ground truth data obtained from Gazebo with the values obtained by solving the NLS problem (32) with the help of (19). To this end, we drive the Husky robot in Gazebo on synthetic terrain models and collect the odometry data, as shown in Fig. 3(a). We then compare the ground truth values $z_{gt}$, $\beta_{gt}$, and $\gamma_{gt}$ from Gazebo with the predicted $z_{pred}$, $\beta_{pred}$, and $\gamma_{pred}$ from our pose optimizer. Several runs of the robot on different terrains were performed, and some qualitative results are shown in Figures 3(b) and 3(c). We can see that the predicted pose values closely match the ground truth values. The statistical results presented in Figures 3(d) and 3(e) show the minimum, maximum, and median error obtained across several runs. It can be seen that our NLS pose optimizer is highly accurate in predicting the pose of the robot on different terrains, and the error values are minimal.
@@ V-C Validation of the Safe Planning on Uneven Terrain
Next, we present a statistical result in Fig. 6(a). Here, we also highlight the significance of the weighting term $w_{\theta}$ in the stability cost 18. Several trajectories were simulated on different terrains, and the average worst-case tip-over angles were recorded. It can be seen that for the trajectories without stability cost, the value is lower. In the case of tip-over angle, higher positive values are preferable, as shown in (15). The values with stability cost are far higher than those without stability cost, which confirms that the trajectories taken by the vehicle with stability cost have a higher safety margin. It can also be seen that as the weight $w_{\theta}$ increased from 0.05 to 0.2, the values of the angles increased, which in turn increased the tip-over stability margin.
@@ V-D Comparison with CEM sampling-based planner
This section investigates the importance of differentiability of the wheel-terrain interaction/pose predictor model. If the implicit gradients are not available, then one can use a black-box sampling-based optimizer such as CEM. However, it will require an order of magnitude larger calls of the NLS pose predictor. In fact, it may not even be possible to solve our bi-level optimizer through a sampling-based approach on resource-constrained hardware. Nevertheless, in this section, our aim is to show that our gradient-based approach is competitive with a more computationally demanding approach. The results are summarized in Fig 6(b), which shows the statistical results of the average worst-case tip-over angle obtained with our gradient-based approach and CEM. Two different batch sizes (100, 20) were used for CEM. The results show that our gradient descent-based planner performs as well as a CEM planner with batch size 100 and is slightly better than a CEM with batch size 20. The average computation time for our approach was $0.42s$, whereas the CEM with batch sizes 100 and 20 took $2.12s$ and $1.03s$, respectively. It is also to be noted that the CEM planners are memory-hungry; thus, implementation on onboard computers might be difficult, if at all possible.
@@
We presented, for the first-time, a differentiable wheel-terrain interaction/pose predictor model that is derived purely from the first principles and yet generalizes to arbitrary terrains. The fidelity of our model was shown to be close to that of a state-of-the-art physics engine. This opens up exciting new possibilities. For example, our pose predictor can be used as a differentiable and parallelizable world model within trajectory optimization or even reinforcement learning pipelines. We presented one such approach based on bi-level optimization, wherein we leveraged implicit gradients from the NLS solver to efficiently compute the optimal trajectory. The proposed optimizer can accommodate arbitrary stability metrics that depend on the vehicle’s pose on the uneven terrain. As an example, we used the force-angle measure and demonstrated stable trajectory planning on uneven terrains. Our future endeavors are focused towards extending the NLS-based approach to handle wheel-terrain dynamics.

### v3

@@ Abstract
Navigation of wheeled vehicles on uneven terrain necessitates going beyond the 2D approaches for trajectory planning. Specifically, it is essential to incorporate the full $6dof$ variation of vehicle pose and its associated stability cost in the planning process. To this end, most recent works aim to learn a neural network model to predict vehicle evolution. However, such approaches are data-intensive and fraught with generalization issues.
@@
We improve the state-of-the-art in the following respects. First, we show that our NLS-based pose prediction closely matches the output of a high-fidelity physics engine. This result, coupled with the fact that we can query gradients of the NLS solver, makes our pose predictor a differentiable wheel-terrain interaction model. We further leverage this differentiability to efficiently solve the proposed bi-level trajectory optimization problem. Finally, we perform extensive experiments and comparisons with a baseline to showcase the effectiveness of our approach in obtaining smooth, stable trajectories.
@@ I Introduction
Applications like forestry, construction, and search and rescue require wheeled vehicles to navigate over uneven terrains. Thus, for automation in these use cases, we need a trajectory planner that can account for the vehicle tip-over stability over uneven terrains. This, in turn, requires explicitly predicting the vehicle’s $6dof$ pose on the terrain and the kinematic and stability costs associated with those poses.
@@
Algorithmic Contribution: We jointly view the wheeled vehicle and the underneath terrain through the lens of closed-loop kinematics, typically used for modeling parallel manipulators. This, in turn, leads to a set of coupled non-linear equations that dictate how the vehicle’s $6dof$ pose depends on the terrain surface geometry. The pose prediction can then be obtained by minimizing the residual of the non-linear equations, which, in our case, is done through a least-squares optimization problem. Importantly, our set-up allows for differentiating through the least squares solver using implicit function theorem to ascertain how the pose of the vehicle will vary if its position on the terrain is perturbed. Thus, in essence, our non-linear least squares (NLS) approach provides a differentiable wheel-terrain interaction model, albeit at the kinematic level.
@@ III-C Tip-over stability criteria
$\boldsymbol{\pi}_{i}$ $=(\mathbf{I}-\mathbf{\hat{e}}_{i}\mathbf{\hat{e}}_{i}^{T})(\mathbf{p}_{oc,i+1}-\mathbf{p}_{og}),$ (9)
$\boldsymbol{\nu}_{i}$ $=(\mathbf{I}-\mathbf{\hat{e}}_{i}\mathbf{\hat{e}}_{i}^{T})\boldsymbol{\nu},$ (10)
where $\mathbf{I}$ is the identity matrix and $\boldsymbol{\nu}$ is the total force acting on the vehicle. The angles between the forces and the tip-over axis normals can be computed as:
@@ IV-A Functional Form for the Terrain Model
{dmath}
f(x^j,y^j) = ∑_n=1^N a_ncos(ω_1,n x^j +ω_2,n y^j ) + b_nsin(ω_3,n x^j +ω_4,n y^j ),
@@ IV-B Differentiable Wheel-Terrain Interaction
We can expand (20) using the Euler-angle parametrization for the rotation matrix $\mathbf{R}$. When the resulting expressions are stacked alongside (24), we obtain a set of 16 coupled non-linear equations in the following form:
@@ IV-C Bi-Level Optimization Based Trajectory Planning
where $\boldsymbol{\xi}=[\mathbf{c}_{x},\mathbf{c}_{y}]$ and $\mathbf{u}^{*}$ is formed by stacking $\mathbf{u}_{k}^{*}$ at different instants of time. The matrices $\mathbf{A}_{eq},\mathbf{A}$ and the vectors $\mathbf{b}_{eq},\mathbf{b}$ comes from (28b)-(28e).
@@ IV-D Projected gradient descent
The optimal value for the parameter $\boldsymbol{\xi}$ is obtained through projected gradient descent. The gradient descent update rule is given as
@@
${}^{t+1}\boldsymbol{\xi}=$ $\arg\min_{\boldsymbol{\xi}}\frac{1}{2}\left\lVert{}^{t+1}\overline{\boldsymbol{\xi}}-^{t+1}\boldsymbol{\xi}\right\rVert_{2}^{2},$ (31)
@@ V-A Implementation Details
The trajectory planner comprising of the bi-level optimization (29a)-(29c) and projection (32)-(33) is implemented in Python using the JAX [23] library for GPU-accelerated computing. The matrix W in (7) is constructed from a $10^{th}$ order polynomial. Our simulation pipeline consists of open-loop analyses carried out with stand-alone Python code, and the closed-loop simulations were performed on the Gazebo simulator [24]. The terrains were created using Blender [25] and were converted to the required formats for the Python and Gazebo simulations. All simulations were conducted for $100$ time-steps with a sampling time of $0.2s$ on a computer with Intel i9 CPU and Nvidea RTX 3080 GPU. For modeling the terrain at each instant of time, we take a patch of radius $7m$ centered at the vehicle position. The number of Fourier frequencies $N$ in the terrain model was taken as $100$ to balance the trade-off between computation time and accuracy.
@@ V-B Validation using the Gazebo simulator
In this section, we analyze the accuracy of our pose predictor by comparing the ground truth data obtained from Gazebo with the values obtained by solving the NLS problem (26) with the help of (17). To this end, we drive the Husky robot in Gazebo on synthetic terrain models and collect the odometry data, as shown in Fig. 4(a). We then compare the ground truth values $z_{gt}$, $\beta_{gt}$, and $\gamma_{gt}$ from Gazebo with the predicted $z_{pred}$, $\beta_{pred}$, and $\gamma_{pred}$ from our pose optimizer. Several runs of the robot on different terrains were performed, and some qualitative results are shown in Figures 4(b) and 4(c). We can see that the predicted pose values closely match the ground truth values. The statistical results presented in Figures 4(d) and 4(e) show the minimum, maximum, and median error obtained across several runs. It can be seen that our NLS pose optimizer is highly accurate in predicting the pose of the robot on different terrains, and the error values are minimal.
@@ V-C Validation of the Safe Planning on Uneven Terrain
Next, we present a statistical result in Fig. 7(a). Here, we also highlight the significance of the weighting term $w_{\theta}$ in the stability cost 16. Several trajectories were simulated on different terrains, and the average worst-case tip-over angles were recorded. It can be seen that for the trajectories without stability cost, the value is lower. In the case of tip-over angle, higher positive values are preferable, as shown in (13). The values with stability cost are far higher than those without stability cost, which confirms that the trajectories taken by the vehicle with stability cost have a higher safety margin. It can also be seen that as the weight $w_{\theta}$ increased from 0.05 to 0.2, the values of the angles increased, which in turn increased the tip-over stability margin.
@@ V-D Comparison with CEM sampling-based planner
This section investigates the importance of differentiability of the wheel-terrain interaction/pose predictor model. If the implicit gradients are not available, then one can use a black-box sampling-based optimizer such as CEM. However, it will require an order of magnitude larger calls of the NLS pose predictor. In fact, it may not even be possible to solve our bi-level optimizer through a sampling-based approach on resource-constrained hardware. Nevertheless, in this section, our aim is to show that our gradient-based approach is competitive with a more computationally demanding approach. The results are summarized in Fig 7(b), which shows the statistical results of the average worst-case tip-over angle obtained with our gradient-based approach and CEM. Two different batch sizes (100 and 20) were used for CEM. The results show that our gradient descent-based planner performs as well as a CEM planner with batch size 100 and is slightly better than a CEM with batch size 20. The average computation time for our approach was $0.42s$, whereas the CEM with batch sizes 100 and 20 took $2.12s$ and $1.03s$, respectively. It is also to be noted that the CEM planners are memory-hungry; thus, implementation on onboard computers might be difficult, if at all possible.
@@
We presented, for the first time, a differentiable wheel-terrain interaction/pose predictor model that is derived purely from the first principles and yet generalizes to arbitrary terrains. The fidelity of our model was shown to be close to that of a state-of-the-art physics engine. This opens up exciting new possibilities. For example, our pose predictor can be used as a differentiable and parallelizable world model within trajectory optimization or even reinforcement learning pipelines. We presented one such approach based on bi-level optimization, wherein we leveraged implicit gradients from the NLS solver to efficiently compute the optimal trajectory. The proposed optimizer can accommodate arbitrary stability metrics that depend on the vehicle’s pose on uneven terrains. As an example, we used the force-angle measure and demonstrated stable trajectory planning on uneven terrains. Our future endeavors are focused on extending the NLS-based approach to handle wheel-terrain dynamics.